2 subtracted from the product of 5 and a number is 73. how would i write that as an equation
Answer by farohw(175)


2 subtracted from the product of 5 and a number is 73.

The product of 5 and a number is 5*x (or any variable like y, z or n, etc. that represents "a number").

Therefore, 5x - 2 = 73 is your equation.

To solve,
5x - 2 = 73
5x = 73 + 2
5x = 75
5x/5 = 75/5
x = 15